# Grover Speedup from Many Forms of the Zeno Effect

### v3

@@ 4.1 Zeno Effect Induced by Projective Measurements
Now that we have discussed the already well-understood case of continuous time quantum walk and adiabatic evolution, let us consider extending the previous work in [46]. As was discussed in that work, a projective measurement between the two lowest eigenstates of a search Hamiltonian can be performed in a time which is proportional to the inverse of the instantaneous gap between the ground and first-excited state of that Hamiltonian. This was argued by considering a model of measurement involving a continuous variable which, when compiled to a gate-model quantum computer becomes the well-known phase estimation algorithm. We discuss the details of this model later in section 5.2 where we use it to derive a physically realistic description of dissipation.
@@
As was observed in [46], if we measure exactly at the avoided crossing where $|g\rangle=\frac{1}{\sqrt{2}}(|\tilde{\omega}\rangle+|m\rangle)$ and $|e\rangle=\frac{1}{\sqrt{2}}(|\tilde{\omega}\rangle-|m\rangle)$, the result of a single measurement is $\frac{1}{2}\left(|g\rangle\!\langle g|+|e\rangle\!\langle e|\right)=\frac{1}{2}\left(|\tilde{\omega}\rangle\!\langle\tilde{\omega}|+|m\rangle\!\langle m|\right)$. Therefore a measurement in this basis will result in a $50$% chance of finding the marked state.
It follows from the structure of the eigenstates that measuring anywhere near the avoided crossing will result in a non-zero probability of $|m\rangle$. Moreover, performing a projective measurement can only decrease the purity of a density matrix. As argued in appendix C the probability of being found in the marked state can be lower bounded in terms of the purity of the density matrix,
@@
As long as at least one measurement is applied in the vicinity of the avoided crossing in a way which decreases the purity significantly, a sequence of measurements defines a valid algorithm which gives a quantum speedup. For the purpose of this paper the most interesting of these are measurements performed at sequentially decreasing values of $f(\tau)$ in a way which manifests as a (discrete) Zeno effect. Performing a large but not strictly infinite number of measurements yields a valid algorithm, however this will not have a well defined continuum limit, since the total runtime will approach infinity. To have a well defined continuum limit, we must define an operation which behaves like a projective measurement when performed for a time proportional to the gap, but is also meaningful for shorter times. We argue in the next section that decoherence in the diagonal basis of the Hamiltonian is one such operation.
An astute reader will notice that the arguments in the last paragraph do not strictly apply to a Zeno effect, since in this case the measurements do not significantly decrease the purity. We can however show numerically that a sequence of measurements can lead to a Zeno effect. Figure 5 illustrates (as filled circles) that indeed if an increasing number of measurements are performed at points determined by the optimal adiabatic schedule the marked state probability tends to unity. Moreover, we argue analytically that a Zeno effect of this form can effectively maintain the ground state in appendix C.
Figure 5: Marked state probability versus number of operations (either dephasing or destruction) performed (with placement determined by the optimal adiabatic schedule as depicted in figure 1) for a variety of operations. Filled circles correspond to performing a number of projective measurements in the instantaneous energy eigenbasis of eq. 2 defined on the x-axis, while filled squares correspond to the same but partially destructive measurements. Unfilled symbols correspond to the partial counterparts (decoherence and dissipation respectively) where the total rotation angles add to the fixed value for a single full operation.
@@ 4.2 Zeno Effect Induced by Decoherence
where $u(\tau)\leq 1$ and this is the term where the decoherence is manifested. Since decoherence decreases $\mathrm{Tr}\left[\rho^{2}\right]$, than by eq. 14, any protocol containing a single application of decoherence in the vicinity of the avoided crossing where $u(t_{\mathrm{scale}})$ becomes significantly less than one will lead to an optimal algorithm.
This is getting ahead of ourselves a bit however, we first need to show that there is a way to apply decoherence in a way which depends only on rescaled time, in other words with a time scale proportional to $g^{-1}_{\mathrm{min}}$. Starting with applications at at fixed $f(\tau)=\gamma$, and then extending to multiple applications of decoherence. There will be multiple ways to show this, but we elect to build on results we already have for quantum walk. Consider a case where instead of just evolving with the Hamiltonian in equation 8, we instead flip a coin, if it is heads we apply
@@
this dependence means that for any value of $\tau$ there is a fixed rotation angle $\phi$, where $\sqrt{\gamma^{2}+1}$ is the rescaled energy gap between the ground and first excited state. In particular total dephasing, with $\phi=\pi/2$ is identical to projective measurement. For simplicity we initially focus on a constant $f(\tau)=\gamma$.
Figure 6: Marked state probability versus number of dephasing operations with different rotation angles per operation (encoded as colour/grayscale) as depicted on the bar). Circles (gold) are used to indicate the projective measurement case. The values of $\tau$ for which the operations are performed are chosen to match the optimal adiabatic schedule.
@@
In other words in the continuum limit, the decoherence vanishes if defined in this way. The approach to this limit can be seen in figure 5 (illustrated by the hollow circles). The reason for the somewhat counter-intuitive behaviour of this limit is actually an unintended Zeno effect, if we consider the effect of reversing the time evolution within the single avoided crossing model, this is equivalent to applying an operation to switch the role of the ground and first excited state, and then applying another one at the end. If we consider a specific set of choices to reverse or not reverse the time evolution at each time step $r_{j}\in\{0,1\}$, we have (building on eq. 9)
@@
$=\tilde{X}^{\sum_{j}r_{j}}\mathcal{T}\prod^{m}_{j=1}\exp\left(\frac{-it_{\mathrm{scale}}\sqrt{\gamma_{j}+1}}{m}\tilde{Z}\right)\tilde{X}^{r_{j}}|\psi(t=0)\rangle,$ (22)
@@
In this case, as long as $m>q$, than on average $q$ flips will occur, but this will not scale with $m$, so therefore we can set $q$ large enough to achieve a good approximation of continuous decoherence, but not strictly send it to infinity as we have to do to $m$ to achieve a continuum limit. The effective coupling rate (for a given $f(\tau)=\gamma$) for this decoherence is
$\kappa\approx\frac{q}{t_{\mathrm{scale}}}(1-\cos\left(\sqrt{f^{2}(\tau)+1}\frac{t_{\mathrm{scale}}}{q}\right))\approx t_{\mathrm{scale}}\frac{f^{2}(\tau)+1}{2q}=\left(\frac{g(\tau)}{g_{\mathrm{min}}}\right)^{2}\kappa_{0},$ (25)
@@
Where we note that the usual time evolution component of the Lindblad equation has vanished due to frequent reversal of the time evolution. This equation will become exact in the limit where $q\rightarrow\infty$, but is only nontrivial if $\kappa$ remains finite, which further implies that $t_{\mathrm{scale}}\propto q$ and therefore also $t_{\mathrm{scale}}\rightarrow\infty$. A more detailed derivation can be found in section D.1. From here the arguments that this decoherence can lead to an optimal algorithm follow directly from the fact that the purity of the density matrix is reduced by the evolution, and therefore, from eq. 14, overlap with state $|m\rangle$ is unavoidable.
As before this argument does not strictly apply to cases where a Zeno effect is manifested, since these systems will remain in an approximately pure state throughout, we examine these cases numerically to show that indeed a Zeno effect can lead to finding the marked state with nearly $100\%$ probability. Figure 7 shows that a continuous Zeno effect does indeed lead to success probability which approaches $100\%$. We also show an alternative model for decoherence in appendix E.
Figure 7: Probability of finding the marked state versus total scaled runtime for a Zeno effect due to continuous dissipation, modeled by equation 28.
@@ 5.1 Zeno Effect from Partially Destructive Measurements
Another kind of Zeno effect can be manifested due to dissipation, somewhat counter-intuitively a stronger dissipation can cause a Zeno effect which prevents the state which is coupled to dissipation from being occupied. As we did in the case of projective measurement and decoherence, we start with a model based on discrete measurements and then perform a continuum extension. To start with, we modify the model of measurement given in equation 13 to consider instead what we call a “partially destructive” measurement. When performing such a measurement, if the system is in $|g\rangle$ than it remains in $|g\rangle$, but if it is in state $|e\rangle$ it is sent into a “destroyed” state $|d\rangle$. The state $|d\rangle$ is considered computationally useless, if the system is in this state at the end of the computation, than the computation has failed and must be re-attempted. Perhaps unsurprisingly we will see that by taking advantage of the Zeno effect protocols can be developed in which the probability of being found in $|d\rangle$ can be made arbitrarily small. In terms of density matrices, this operation takes the form
@@
The algorithm given in [46] will also work for this kind of measurement. A partially destructive measurement of this type right at the avoided crossing would result in the $|d\rangle$ state $50\%$ of the time, but would be found in the $|g\rangle$ state the other $50\%$, and of these cases a subsequent measurement would find $|m\rangle$ $50\%$ of the time, for an overall success probability of $25\%$ each time the algorithm is run, since this is still $O(1)$ then such an algorithm would still yield an optimal speedup. Similar arguments apply for measurements anywhere close to the avoided crossing. However, since amplitude in the excited state is “lost” this case does not have the elegant argument about purity which could be made in the projective case. In particular, a pair of sequential measurements where $|\gamma|$ takes the same large value, but has the opposite sign will result in the system being found completely in the $|d\rangle$ state since $|e\rangle$ with $\gamma$ would be $|g\rangle$ with $-\gamma$ and vice-versa.
Zeno effects are however still possible, and indeed as the number of measurements are increased can approach $100\%$ success as can be seen in figure 8. As with the case of projective measurement, there is no way to directly take a continuum limit of the partially destructive measurements, we instead need to take a continuous quantum channel which behaves like a partially destructive measurement, which as we discuss in the next section is dissipation out of the $|e\rangle$ state.
@@ 5.2 Zeno Effect from Dissipation
The action of this Hamiltonian evolution is to send $|E_{a}\rangle|x=0\rangle$ to $|E_{a}\rangle\Bigl{\lvert}x=tE_{a}\Bigr{\rangle}$, wavepackets move distances apart corresponding to the energy differences multiplied by time. As was discussed in [46] this model can be compiled to a universal quantum computer, and in fact corresponds to the well known phase estimation algorithm in that case. In light of this fact, it is clear that this model can be considered physically realistic.
@@
where we note that all variables except for the original system variable are returned to their original state for the next application of dissipation. The model we have shown here can also be straightforwardly adapted to describe decoherence giving an alternative to the model in section 4.2, we show this construction explicitly in appendix E.
@@
Equipped with this model, we can now perform multiple dissipative steps with a scaled rotation angle $\phi$, following equation 19, and setting $t_{\mathrm{rot}}=\pi/2$. The results can be found in figure 8 and indeed show an approach to the behaviour of repeated partially destructive measurements. Before discussing the continuum limit it is worth discussing a subtlety of this model.
@@
Naively, from the equations we have written previously, it would appear that this dissipation model could yield a speedup which is even faster than the $\sqrt{N}$ Grover speedup. Namely, if we were able to take $t_{\mathrm{rot}}\propto N^{\frac{1}{4}}$ and $t_{\mathrm{scale}}\propto N^{\frac{1}{4}}$, we could obtain the same dynamics, but in a time which scales only as $N^{\frac{1}{4}}$. The reason that such an implementation is unphysical (it has to be since it gives a scaling which is not possible in real quantum systems) is somewhat subtle. If we scaled in this way $x_{mag}$ scales as $N^{-\frac{1}{4}}$. In such a scenario, the width of the initial state of the continuous variable which we have assumed are “sufficiently narrow” to be approximated as classical particles, would have to also scale down at least as $N^{-\frac{1}{4}}$, such scaling is physically unrealistic, and without it, the final wavepackets would increasingly overlap as $N$ scales, for a large $N$ it would not be possible to reliably distinguish $|e\rangle$ from $|g\rangle$.
In the actual model we have considered, the final difference in the continuum variable values does not scale with $N$, so therefore the same “sufficiently narrow” width could work for all $N$ values. Similar arguments follow for all aspects of the interaction between the additional qubit and the variable, since all of the scaling is confined between the interaction between the single avoided crossing system and the continuum variable.
We now consider what happens in the continuum limit. As with the dissipation case, we can see that as $t_{\mathrm{scale}}$ is decreased the probability of measuring in $|d\rangle$ after applying $D(\tau,t_{\mathrm{scale}}/q,t_{\mathrm{rot}})$ will scale as $1-\sin^{2}\left(2t_{\mathrm{rot}}x_{\mathrm{mag}}\right)\propto t^{2}_{\mathrm{scale}}/q^{2}$. Therefore if we divide the evolution into $q$ segments with $q\rightarrow\infty$, no dynamics will happen unless we also scale $t_{\mathrm{scale}}\propto q$. The reason for this is actually another Zeno effect, we are measuring the rotated qubit with increasing frequency, eventually stopping its dynamics. This behaviour is illustrated numerically in figure 5 as hollow squares.
@@
In the limit described above, infinite $m$ and large but finite $q$, each stage measurement effectively describes a fraction of the probability to be in state $|e\rangle$ being moved to $|d\rangle$ (represented by aborting the computation in the event of a $|1\rangle$ measurement). For a single measurement the probability of this happening is $\sin^{2}\left(2t_{\mathrm{rot}}x_{\mathrm{mag}}\right)$ according to equation 34. For large $q$ this can be converted into a decay rate by dividing by time
@@
as before we can define $\kappa_{0}$ in terms of the Zeno time from equation 27,
@@
Since the action of sending $|e\rangle\rightarrow|d\rangle$ can be represented as a Lindblad master equation with a Lindblad operator of $L=|d\rangle\!\langle e|$, the overall differential equation in the continuum limit becomes (see derivation in appendix D.2 )
@@
Figure 9: Probability of finding the marked state versus total scaled runtime for a Zeno effect due to continuous dissipation, modeled by equation 40.
It is worth remarking that the same steps could be followed, but where the measured qubit is reset to $|0\rangle$ and the computation continued rather than aborting upon measuring $|1\rangle$ this would correspond to a Zeno effect through weak measurement which would have identical action to the decoherence examined previously, this is demonstrated explicitly in appendix E.

### v4

@@ 4.1 Zeno Effect Induced by Projective Measurements
Now that we have discussed the already well-understood case of continuous-time quantum walk and adiabatic evolution, let us consider extending the previous work in [47]. As was discussed in that work, a projective measurement between the two lowest eigenstates of a search Hamiltonian can be performed in a time which is proportional to the inverse of the instantaneous gap between the ground and first-excited state of that Hamiltonian. This was argued by considering a model of measurement involving a continuous variable which, when compiled to a gate-model quantum computer, becomes the well-known phase estimation algorithm. We discuss the details of this model later in section 5.2 where we use it to derive a physically realistic description of dissipation.
@@
As was observed in [47], if we measure exactly at the avoided crossing where $|g\rangle=\frac{1}{\sqrt{2}}(|\tilde{\omega}\rangle+|m\rangle)$ and $|e\rangle=\frac{1}{\sqrt{2}}(|\tilde{\omega}\rangle-|m\rangle)$, the result of a single measurement is $\frac{1}{2}\left(|g\rangle\!\langle g|+|e\rangle\!\langle e|\right)=\frac{1}{2}\left(|\tilde{\omega}\rangle\!\langle\tilde{\omega}|+|m\rangle\!\langle m|\right)$. Therefore, a measurement in this basis will result in a $50$% chance of finding the marked state.
It follows from the structure of the eigenstates that measuring anywhere near the avoided crossing will result in a non-zero probability of $|m\rangle$. Moreover, performing a projective measurement can only decrease the purity of a density matrix. As argued in appendix D the probability of being found in the marked state can be lower bounded in terms of the purity of the density matrix,
@@
As long as at least one measurement is applied in the vicinity of the avoided crossing in a way which decreases the purity significantly, a sequence of measurements defines a valid algorithm which gives a quantum speedup. For the purpose of this paper, the most interesting of these are measurements performed at sequentially decreasing values of $f(\tau)$ in a way which manifests as a (discrete) Zeno effect. Performing a large but not strictly infinite number of measurements yields a valid algorithm, however this will not have a well-defined continuum limit, since the total runtime will approach infinity. To have a well-defined continuum limit, we must define an operation which behaves like a projective measurement when performed for a time proportional to the gap, but is also meaningful for shorter times. We argue in the next section that decoherence in the diagonal basis of the Hamiltonian is one such operation.
An astute reader will notice that the arguments in the last paragraph do not strictly apply to a Zeno effect, since in this case the measurements do not significantly decrease the purity. We can however show numerically that a sequence of measurements can lead to a Zeno effect. Figure 5 illustrates (as filled circles) that indeed if an increasing number of measurements are performed at points determined by the optimal adiabatic schedule, the marked state probability tends to unity. Moreover, we argue analytically that a Zeno effect of this form can effectively maintain the ground state in appendix D.
Figure 5: Marked state probability versus number of operations (either dephasing or destruction) performed (with placement determined by the optimal adiabatic schedule, as depicted in figure 1) for a variety of operations. Filled circles correspond to performing a number of projective measurements in the instantaneous energy eigenbasis of eq. 2 defined on the x-axis, while filled squares correspond to the same but partially destructive measurements. Unfilled symbols correspond to models (based on decoherence and dissipation, respectively) where the total rotation angles add to the fixed value for a single full operation. The filled symbols on the other hand correspond to versions where each operation corresponds to full dephasing (in the case of decoherence) or full removal of amplitude in the excited state (in the case of dissipation).
@@ 4.2 Zeno Effect Induced by Decoherence
where $u(\tau)\leq 1$ and this is the term where the decoherence is manifested. Since decoherence decreases $\mathrm{Tr}\left[\rho^{2}\right]$, then by eq. 14, any protocol containing a single application of decoherence in the vicinity of the avoided crossing where $u(t_{\mathrm{scale}})$ becomes significantly less than one will lead to an optimal algorithm.
This is getting ahead of ourselves a bit however, we first need to show that we can apply decoherence in a way which depends only on rescaled time, in other words with a timescale proportional to $g^{-1}_{\mathrm{min}}$. Starting with applications at fixed $f(\tau)=\gamma$, and then extending to multiple applications of decoherence. There will be multiple ways to show this, but we elect to build on results we already have for quantum walk. Consider a case where instead of just evolving with the Hamiltonian in equation 8, we instead flip a coin, if it is heads we apply
@@
this dependence means that for any value of $\tau$ there is a fixed rotation angle $\phi$, where $\sqrt{\gamma^{2}+1}$ is the rescaled energy gap between the ground and first excited state. In particular, total dephasing, with $\phi=\pi/2$ is identical to projective measurement. For simplicity, we initially focus on a constant $f(\tau)=\gamma$.
Figure 6: Marked state probability versus number of dephasing operations with different rotation angles per operation (encoded as colour/grayscale as depicted on the bar). Circles (gold) are used to indicate the projective measurement case. The values of $\tau$ for which the operations are performed are chosen to match the optimal adiabatic schedule.
@@
In other words, in the continuum limit, the decoherence vanishes if defined in this way. The approach to this limit can be seen in figure 5 (illustrated by the hollow circles). The reason for the somewhat counter-intuitive behaviour of this limit is actually an unintended Zeno effect, if we consider the effect of reversing the time evolution within the single avoided crossing model, this is equivalent to applying an operation to switch the role of the ground and first excited state, and then applying another one at the end. If we consider a specific set of choices to reverse or not reverse the time evolution at each time step $r_{j}\in\{0,1\}$, we have (building on eq. 9)
@@
$=\tilde{X}^{\sum_{j}r_{j}}\mathcal{T}\prod^{m}_{j=1}\exp\left(\frac{-it_{\mathrm{scale}}\sqrt{\gamma^{2}_{j}+1}}{m}\tilde{Z}\right)\tilde{X}^{r_{j}}|\psi(t=0)\rangle,$ (22)
@@
In this case, as long as $m>q$, then on average $q$ flips will occur, but this will not scale with $m$, so therefore we can set $q$ large enough to achieve a good approximation of continuous decoherence, but not strictly send it to infinity as we have to do to $m$ to achieve a continuum limit.
One way to define an effective coupling rate is to examine how much, on average, the state probability will transfer from the positive superposition state $(|e\rangle+|g\rangle)/\sqrt{2}$ to the negative superposition state $(|e\rangle-|g\rangle)/\sqrt{2}$ before the transfer direction is reversed. This expression divided by the time taken $\Delta p_{\pm}/\Delta t$ will give a quantity which can be roughly understood as “effective” coupling rate. This is coupling rate in the sense that the continuous process obtained by taking the limit $q\rightarrow\infty$ will effectively have a coupling rate equal to this quantity. Equation 22 implies that the effective coupling rate defined in this way (for a given $f(\tau)=\gamma$) is
$\kappa=\frac{\Delta p_{\pm}}{\Delta t}\approx\frac{q}{t_{\mathrm{scale}}}\left|\frac{\langle e|-\langle g|}{\sqrt{2}}\exp\left(\frac{-it_{\mathrm{scale}}\sqrt{f^{2}(\tau)+1}}{m}\tilde{Z}\right)\frac{|e\rangle+|g\rangle}{\sqrt{2}}\right|^{2}$
$=\frac{q}{2t_{\mathrm{scale}}}(1-\cos\left(\sqrt{f^{2}(\tau)+1}\frac{2t_{\mathrm{scale}}}{q}\right))\approx t_{\mathrm{scale}}\frac{f^{2}(\tau)+1}{2q}=\left(\frac{g(\tau)}{g_{\mathrm{min}}}\right)^{2}\kappa_{0},$ (25)
@@
Where we note that the usual time evolution component of the Lindblad equation has vanished due to frequent reversal of the time evolution. This equation will become exact in the limit where $q\rightarrow\infty$, but is only nontrivial if $\kappa$ remains finite, which further implies that $t_{\mathrm{scale}}\propto q$ and therefore also $t_{\mathrm{scale}}\rightarrow\infty$. A more detailed derivation can be found in section E.1. From here the arguments that this decoherence can lead to an optimal algorithm follow directly from the fact that the purity of the density matrix is reduced by the evolution, and therefore, from eq. 14, overlap with state $|m\rangle$ is unavoidable.
As before this argument does not strictly apply to cases where a Zeno effect is manifested, since these systems will remain in an approximately pure state throughout, we examine these cases numerically to show that indeed a Zeno effect can lead to finding the marked state with nearly $100\%$ probability. Figure 7 shows that a continuous Zeno effect does indeed lead to success probability which approaches $100\%$. We also show an alternative model for decoherence in appendix F.
Figure 7: Probability of finding the marked state versus total scaled runtime for a Zeno effect due to continuous dissipation, modelled by equation 28.
@@ 5.1 Zeno Effect from Partially Destructive Measurements
Another kind of Zeno effect can be manifested due to dissipation, somewhat counter-intuitively a stronger dissipation can cause a Zeno effect which prevents the state which is coupled to dissipation from being occupied. As we did in the case of projective measurement and decoherence, we start with a model based on discrete measurements and then perform a continuum extension. To start with, we modify the model of measurement given in equation 13 to consider instead what we call a “partially destructive” measurement. When performing such a measurement, if the system is in $|g\rangle$ than it remains in $|g\rangle$, but if it is in state $|e\rangle$ it is sent into a “destroyed” state $|d\rangle$. The state $|d\rangle$ is considered computationally useless, if the system is in this state at the end of the computation, then the computation has failed and must be re-attempted. Perhaps unsurprisingly we will see that by taking advantage of the Zeno effect, protocols can be developed in which the probability of being found in $|d\rangle$ can be made arbitrarily small. In terms of density matrices, this operation takes the form
@@
The algorithm given in [47] will also work for this kind of measurement. A partially destructive measurement of this type right at the avoided crossing would result in the $|d\rangle$ state $50\%$ of the time, but would be found in the $|g\rangle$ state the other $50\%$, and of these cases a subsequent measurement would find $|m\rangle$ $50\%$ of the time, for an overall success probability of $25\%$ each time the algorithm is run, since this is still $O(1)$ then such an algorithm would still yield an optimal speedup. Similar arguments apply for measurements anywhere close to the avoided crossing. However, since amplitude in the excited state is “lost” this case does not have the elegant argument about purity which could be made in the projective case. In particular, a pair of sequential measurements where $|\gamma|$ takes the same large value, but has the opposite sign will result in the system being found completely in the $|d\rangle$ state since $|e\rangle$ with $\gamma$ would be $|g\rangle$ with $-\gamma$ and vice versa.
Zeno effects are however still possible, and indeed as the number of measurements are increased can approach $100\%$ success, as can be seen in figure 8. As with the case of projective measurement, there is no way to directly take a continuum limit of the partially destructive measurements, we instead need to take a continuous quantum channel which behaves like a partially destructive measurement, which as we discuss in the next section is dissipation out of the $|e\rangle$ state.
@@ 5.2 Zeno Effect from Dissipation
The action of this Hamiltonian evolution is to send $|E_{a}\rangle|x=0\rangle$ to $|E_{a}\rangle\Bigl{\lvert}x=tE_{a}\Bigr{\rangle}$, wavepackets move distances apart corresponding to the energy differences multiplied by time. As was discussed in [47] this model can be compiled to a universal quantum computer, and in fact corresponds to the well known phase estimation algorithm in that case. In light of this fact, it is clear that this model can be considered physically realistic. A review of this model can be found in appendix A.
@@
where we note that all variables except for the original system variable are returned to their original state for the next application of dissipation. The model we have shown here can also be straightforwardly adapted to describe decoherence giving an alternative to the model in section 4.2, we show this construction explicitly in appendix F.
@@
Equipped with this model, we can now perform multiple dissipative steps with a scaled rotation angle $\phi$, following equation 19, and setting $t_{\mathrm{rot}}=\pi/2$. The results can be found in figure 8 and indeed show an approach to the behaviour of repeated partially destructive measurements. Before discussing the continuum limit, it is worth discussing a subtlety of this model.
@@
Naively, from the equations we have written previously, it would appear that this dissipation model could yield a speedup which is even faster than the $\sqrt{N}$ Grover speedup. Namely, if we were able to take $t_{\mathrm{rot}}\propto N^{\frac{1}{4}}$ and $t_{\mathrm{scale}}\propto N^{\frac{1}{4}}$, we could obtain the same dynamics, but in a time which scales only as $N^{\frac{1}{4}}$. The reason that such an implementation is unphysical (it has to be, since it gives a scaling which is not possible in real quantum systems) is somewhat subtle. If we scaled in this way, $x_{mag}$ scales as $N^{-\frac{1}{4}}$. In such a scenario, the width of the initial state of the continuous variable which we have assumed are “sufficiently narrow” to be approximated as classical particles, would have to also scale down at least as $N^{-\frac{1}{4}}$, such scaling is physically unrealistic, and without it, the final wavepackets would increasingly overlap as $N$ scales, for a large $N$ it would not be possible to reliably distinguish $|e\rangle$ from $|g\rangle$.
In the actual model we have considered, the final difference in the continuum variable values does not scale with $N$, therefore the same “sufficiently narrow” width could work for all $N$ values. Similar arguments follow for all aspects of the interaction between the additional qubit and the variable, since all the scaling is confined between the interaction between the single avoided crossing system and the continuum variable.
We now consider what happens in the continuum limit. As with the decoherence case, we can see that as $t_{\mathrm{scale}}$ is decreased the probability of measuring in $|d\rangle$ after applying $D(\tau,t_{\mathrm{scale}}/q,t_{\mathrm{rot}})$ will scale as $1-\sin^{2}\left(2t_{\mathrm{rot}}x_{\mathrm{mag}}\right)\propto t^{2}_{\mathrm{scale}}/q^{2}$. Therefore, if we divide the evolution into $q$ segments with $q\rightarrow\infty$, no dynamics will happen unless we also scale $t_{\mathrm{scale}}\propto q$. The reason for this is actually another Zeno effect, we are measuring the rotated qubit with increasing frequency, eventually stopping its dynamics. This behaviour is illustrated numerically in figure 5 as hollow squares.
@@
In the limit described above, infinite $m$ and large but finite $q$, each stage measurement effectively describes a fraction of the probability to be in state $|e\rangle$ being moved to $|d\rangle$ (represented by aborting the computation in the event of a $|1\rangle$ measurement). For a single measurement, the probability of this happening is $\sin^{2}\left(2t_{\mathrm{rot}}x_{\mathrm{mag}}\right)$ according to equation 34. For large $q$ this can be converted into a decay rate by dividing by time
@@
as before, we can define $\kappa_{0}$ in terms of the Zeno time from equation 27,
@@
Since the action of sending $|e\rangle\rightarrow|d\rangle$ can be represented as a Lindblad master equation with a Lindblad operator of $L=|d\rangle\!\langle e|$, the overall differential equation in the continuum limit becomes (see derivation in appendix E.2 )
@@
Figure 9: Probability of finding the marked state versus total scaled runtime for a Zeno effect due to continuous dissipation, modelled by equation 40.
It is worth remarking that the same steps could be followed, but where the measured qubit is reset to $|0\rangle$ and the computation continued rather than aborting upon measuring $|1\rangle$ this would correspond to a Zeno effect through weak measurement which would have identical action to the decoherence examined previously, this is demonstrated explicitly in appendix F.